Problem Solving: Solve the problem below. Show your solution.

1. Two protons in an atom are 1.75 m x 10 -8 m from each other. Calculate the electric force between them.

2. Calculate the magnitude of the electrostatic force between a + 6.0µC charged particle and a + 8.0µC charged particle separated by a distance of 0.7 cm.

3. Two charges attract each other with a force of 90 N. How far a 5.0 x 10-6 C charge from a -3.0 x 10 -5 C charge?


1
Expert's answer
2021-04-28T20:19:01-0400

1. The force is


"F=\frac{1}{4\pi\epsilon_0}·\frac{e·e}{r^2}=7.53·10^{-13}\text{ N}."

2. The force is


"F=\frac{1}{4\pi\epsilon_0}·\frac{q_1q_2}{r^2}=8800\text{ N}."

3. The distance can be found from the equation of the force:


"d=\sqrt{\frac{q_1q_2}{4\pi\epsilon_0F}}=0.12\text{ m}."